# An Ontology-Driven Graph RAG for Legal Norms: A Hierarchical, Temporal, and Deterministic Approach

### v4

@@ 4.4 A unified execution strategy
To support the three query patterns exemplified above (point-in-time retrieval; hierarchical impact analysis; provenance reconstruction) we implement a single, planner-guided execution strategy that is both deterministic and configurable. The strategy centralizes query interpretation and then applies a small set of composable steps:
@@
In our reference implementation we use: embeddings with dimension 256 (Qwen-3), cosine similarity with $k=8$ (configurable), and the SnapshotLast temporal policy; all answers explicitly state the temporal/membership policies used and attach a JSON provenance annex when Actions are part of the result. These defaults are configurable and should be reported alongside query results to preserve determinism and user trust.
@@
By centralizing planning and exposing policy choices, the system achieves (i) deterministic point-in-time answers, (ii) scalable hierarchical aggregations, and (iii) auditable provenance reconstruction — all from the same core pipeline. This design keeps the system predictable for legal usages while remaining extensible to other jurisdictions and corpus types.
@@ 5 Discussion
The framework’s determinism amplifies the advantages of structured data but also propagates upstream errors more visibly. Erroneous validity intervals or incorrect Action metadata will lead to systematically flawed retrievals. Mitigation is therefore crucial and relies on practices such as multi-source corroboration, attaching confidence scores to all outputs, and implementing robust validation and rollback workflows.
Scope, Domain Transfer, and Limitations
The proposed ontology is optimally suited to statutory corpora where hierarchy and amendment actions are explicit. Adapting the model to other legal domains requires careful ontological and procedural adjustments. For case law, the focus would shift from textual diffs to modeling citation graphs and doctrinal drift. For contracts or regulations, their structural heterogeneity may demand bespoke segmentation rules. Explicit policies for handling multilingual corpora and resolving complex cross-references are also necessary extensions for a production-ready system.
Generalization to Non-Legal Domains
While developed for law, the framework’s core principles—separating abstract structure from dated textual expressions and modeling change via explicit events—are highly generalizable. The model can be readily adapted to other domains characterized by structured, versioned documents, such as commercial contracts evolving through amendments, or technical documentation tracking software releases. This suggests a broader applicability of our event-centric, policy-driven versioning model for any corpus where structural integrity and historical accuracy are paramount.
@@
To rigorously quantify the framework’s benefits, we advocate for the creation of jurisdictional, ground-truth benchmarks encoding known historical changes. Evaluation must go beyond standard retrieval metrics to include temporal precision/recall, action-attribution accuracy, and provenance completeness. Crucially, user-centred measures, such as time-to-answer in realistic legal research tasks, are needed to validate the practical utility of the approach.
@@
For practitioners, this approach unlocks faster, auditable historical research and deterministic summaries of legislative change. However, we advise conservative operational policies. Outputs should be presented as verified-assistant results, not definitive legal advice, with full provenance exposed for independent verification. Key ethical considerations include transparency in the system’s resolution policies, ensuring equitable access across jurisdictions, and maintaining immutable audit logs for all data curation and graph updates.
@@ 6 Conclusion
This paper introduced an ontology-driven Graph RAG architecture designed to overcome the critical limitations of standard retrieval systems when applied to diachronic legal corpora. Grounded in a formal, LRMoo-inspired model, our framework treats temporal versions, structural components, and legislative Actions as first-class, interconnected entities. This enables a unified, planner-guided query strategy that transforms legal information retrieval from a probabilistic search into a deterministic, auditable process.
@@
  • A formal, multi-layered representation that separates abstract legal Works from their concrete Temporal and Language Expressions, enabling precise, point-in-time reconstruction.
@@
  • The explicit modelling of legislative Actions, which reifies the causes of textual change, making them retrievable, attributable, and assemblable into auditable provenance chains.
  • A unified, policy-driven query framework that leverages the graph structure to deterministically resolve complex requests for point-in-time retrieval, hierarchical impact analysis, and causal-lineage reconstruction.
While the approach requires a principled investment in data curation and governance, it lays a necessary foundation for the next generation of trustworthy legal AI. By making temporality, structure, and causality explicit, our framework offers a practical and robust path toward building AI systems that can reliably support high-stakes legal decision-making.
@@
We identify several promising research directions, including the publication of a formal OWL ontology for the model, the development of annotated benchmarks for reproducible evaluation, the extension of the framework to other legal domains like case law, and the creation of interactive tooling for provenance visualization and legal workflow integration.

### v5

@@ 4.4 A unified execution strategy
To support the three query patterns exemplified above (point-in-time retrieval; hierarchical impact analysis; provenance reconstruction) we propose a single, planner-guided execution strategy that is both deterministic and configurable. The strategy centralizes query interpretation and then applies a small set of composable steps:
@@
A reference implementation of this framework should define and report its operational defaults to ensure reproducibility. For instance, it could specify parameters such as the embedding model and dimension (e.g., Qwen-3 with 256 dimensions), the similarity function and retrieval depth (e.g., cosine similarity with $k=8$), and the default temporal policy (e.g., $\operatorname{SnapshotLast}$). Fundamentally, all answers generated by the system should explicitly state the temporal and membership policies used for that specific query and attach a machine-readable JSON provenance annex when Action nodes are part of the result. These defaults should be configurable, and their disclosure is essential to preserve determinism and user trust.
@@
By centralizing planning and exposing policy choices, the proposed design makes it possible to achieve (i) deterministic point-in-time answers, (ii) scalable hierarchical aggregations, and (iii) auditable provenance reconstruction—all from the same core pipeline. This design keeps the system predictable for legal usages while remaining extensible to other jurisdictions and corpus types.
@@ 5 Discussion
The framework’s determinism amplifies the advantages of structured data but also propagates upstream errors more visibly. Erroneous validity intervals or incorrect Action metadata will lead to systematically flawed retrievals. Mitigation is therefore essential and relies on practices such as multi-source corroboration, attaching confidence scores to all outputs, and implementing robust validation and rollback workflows.
Scope, Adaptability, and Generalization
The proposed ontology is optimally suited to statutory corpora where hierarchy and amendment actions are explicit and canonical. Adapting this model to other domains, whether legal or non-legal, requires careful ontological and procedural adjustments, primarily centered on how "structure" is defined and "change" is tracked.
For other legal domains:
  • Case Law: The focus would shift from textual diffs to modeling citation graphs, judicial hierarchies, and the temporal evolution of legal precedents (doctrinal drift).
  • Contracts/Regulations (within law): Their structural heterogeneity may demand bespoke semantic segmentation rules for graph construction, as their internal organization can be less standardized than statutory law.
The framework’s core principles—separating abstract structure from dated textual expressions and modeling change via explicit events—are, however, highly generalizable beyond the legal sphere. For non-legal domains:
  • Commercial Contracts: These documents often evolve through amendments and restatements, making our versioning model highly applicable.
  • Technical Documentation: Tracking software releases, API changes, and documentation versions (e.g., user manuals, specifications) also aligns well with our event-centric, policy-driven approach.
@@
A rigorous quantitative validation of our framework’s claims of determinism and accuracy requires a dedicated benchmark, which is currently a notable gap in the legal AI community. To address this, we propose the creation of a ground-truth testbed built upon a legislative corpus with a known, complex amendment history (such as the one used in our case study). This testbed would consist of a curated set of query-answer pairs designed to evaluate the core capabilities of any structure- and temporally-aware retrieval system. To support these claims, future work should focus on the following quantitative metrics:
  • For Deterministic Temporal Retrieval (Query Pattern 1):
    • Temporal Precision: The fraction of retrieved Component Temporal Versions (CTVs) that are correctly valid for the target date or interval specified in the query.
    • Temporal Recall: The fraction of all ground-truth relevant CTVs for a given query that the system successfully retrieved.
  • For Hierarchical Impact Analysis (Query Pattern 2):
    • Action-Attribution Accuracy: The F1-score for correctly identifying all legislative Action nodes that acted upon the temporal versions (CTVs) of components within the specified hierarchical scope and time window.
    • Summary Completeness: The percentage of ground-truth changes within the scope that are correctly included in the final generated summary.
  • For Provenance Reconstruction (Query Pattern 3):
    • Causal-Chain Completeness: The percentage of the ground-truth causal chain (the sequence of Action nodes) that was fully and correctly reconstructed by the system.
  • User-Centred Measures: Beyond automated metrics, user studies are essential to measure the practical impact, such as the reduction in time-to-answer for complex legal research tasks and the perceived increase in the trustworthiness and auditability of the generated answers.
We advocate for the development and public release of such an annotated benchmark to enable reproducible evaluation and a fair comparison between different temporally-aware legal retrieval systems.
@@
For practitioners, this approach unlocks faster, auditable historical research and deterministic summaries of legislative change. However, we advise conservative operational policies. The framework is designed to produce verifiable assistant outputs, not definitive legal advice. Its primary role is to accelerate human expertise by providing candidate answers with full provenance exposed for independent verification. This responsible deployment model is guided by several key ethical considerations:
  • Policy Transparency: The system must be transparent about its internal resolution policies (e.g., temporal and scope-membership rules). Users must understand how an answer was constructed to properly assess its context and reliability.
  • Data Equity and Access: The framework’s effectiveness is contingent on the availability of high-quality, machine-readable legal data. A significant ethical risk is creating a "justice gap," where only well-funded jurisdictions can build such systems, potentially marginalizing those with less accessible legal corpora.
  • Curation and Auditability: Maintaining immutable audit logs for all data curation and graph updates is essential. The process of modeling legal history must itself be transparent and traceable to prevent manipulation or the introduction of systemic errors.
@@ 6 Conclusion
This paper introduced the Structure-Aware Temporal Graph RAG (SAT-Graph RAG), an ontology-driven architecture designed to overcome the critical limitations of standard retrieval systems when applied to diachronic legal corpora. Grounded in a formal, LRMoo-inspired model, our framework treats a document’s formal structure, its temporal versions, and the legislative Actions that drive its evolution as first-class, interconnected entities. This enables a unified, planner-guided query strategy that transforms legal information retrieval from a probabilistic search into a deterministic, auditable process.
@@
  • The application of a formal, multi-layered representation, based on [2], that separates abstract legal Works from their concrete Temporal and Language Expressions, enabling precise, point-in-time reconstruction.
@@
  • The reification of legislative Actions as first-class, retrievable entities, which makes the causes of textual change directly queryable, attributable, and assemblable into auditable provenance chains, building upon the event-centric model from [2].
  • A multi-aspect retrieval strategy that textualizes structured metadata and causal events into distinct Text Units, making a norm’s context—not just its content—a first-class, searchable facet.
  • A unified, policy-driven query framework that leverages the graph’s explicit formal structure to deterministically resolve complex requests for point-in-time retrieval, hierarchical impact analysis, and causal-lineage reconstruction.
While the approach requires a principled investment in data curation and governance, it lays a necessary foundation for the next generation of trustworthy legal AI. By making awareness of formal document structure a core principle alongside explicit models of temporality and causality, our framework offers a practical and robust path toward building AI systems that can reliably support high-stakes legal decision-making.
@@
We identify several promising research directions, including the publication of a formal OWL ontology for the model and the development of annotated benchmarks for reproducible evaluation. Future work will also focus on extending the framework by adapting it to other domains and investigating hybrid models that combine our top-down, structural approach with bottom-up, content-driven entity graphs, as described in the original Graph RAG proposal [3]. Finally, we envision the creation of interactive tooling for provenance visualization and legal workflow integration.